24 June 2014 UT DDT night

Summary

Conditions were very good. There was some smoke to the SW from a fire in the Galiuros, and we could smell it at sunset, but not later. Seeing was stable at sub-arcsecond for most of the night - only at the end, on the last field which was in the SW, did it start to vary and go up to 1.5". We started with the LBCs and observed the Saturn field. Red IQ was good, Blue IQ was its usual self; but it started out well and seemed to degrade after dofpia. Next switched to MODS and observed PDS456 and 2 of the fainter quasars from the CIV set since conditions were good. We observed V 338 del and then switched back to the LBCs. Seeing continued to be sub-arcsecond on the first field, but on the search field, it started to vary significantly and the backgrounds rose from image to image as the field was setting. Observed the field twice before a significant rise in background counts was seen.

We obtained LBC flat fields at g' and r' at the start and end of the night, and a set of MODS calibrations at the end of the night.

I started the night without updating the UT in the MODS gui, so the mods data will have the 20140623 UT in the filename.

Plan

LBC (1 hr): 21:15-22:15 MST (twilight ends 21:13)
  • 2 x OB20 (a 5-pt dither of 240-sec exposures): dofpia, OB20, dofpia (if needed), OB20

MODS: switch at 22:15 MST. 22:45-01:50 MST
  • qsos: 22:45 - 01:30 MST
    • 22:45-23:08 (PA coincides well enough with PARANGLE until 23:08 MST)
      • start immediately as soon as possible:
        • acqMODS PDS456clumps.acq (22:50)
        • execMODS PDS456_600sec.obs (2 x 600s... 23:15)
      • obtain 15-sec i-band image afterwards (...23:30 MST)
    • 23:30 - 00:15 Q1614+051 (V=19.5)
      • acqMODS Q1614+051.acq (PA=25 for HA=0 ...2.5)
      • execMODS Q1614+051.obs (3x600sec).
      • slit image (imaging, red imcslock, roi 1K x 1K, exptime = 60).
    • 00:15-00:40 BD33 (5" 15 min and 1" 11 min) 25 minutes
    • good conditions:
      • 00:40-01:10 Q1705+018 V=18.9 (PA=30 for HA=1...2) 3x300sec
      • 01:10-01:45(01:55)... Q1702+298 V=19.1 (PA=75 for HA=0 ... 3) 3x400 sec
    • poor conditions
      • Q1656+477 3 x 300 sec
      • Q1658+575 3 x 300 sec
      • Q1816+475 3 x 300 sec
      • Q1857+566 3 x 300 sec
      • Q2048+196 3 x 300 sec
  • 01:45 - 02:15 V335del (8 min acq + 5x60= ~20 minutes)

LBC (1.2 hr): switch at 02:00 MST (observe from 02:20 - 03:31; twilight begins at 03:31 MST)
  • OB21 (a 5-pt dither of 60-sec exposures): dofpia, OB21
  • 2 x OB22 (a 5-pt dither of 240-sec exposures): dofpia, OB22, dofpia (if needed), OB22

Log

Opened after sunset. Clear skies, however, there is a smoke plume from a fire to the SW. The plume extends to the NW and there is a definite smell of smoke in the air. There's a fire in the Galiueros (3000-4000 acres, but the crew can see back side of Mt. Graham, and all OK. But smoke may be coming our way.

03:07 UT Sky flats at g and r at PA=0. Starting at 20k counts...down to 10k.

03:11 Sky flats at g and r at PA=180 (scale x2.5). Starting at 14k,10k... down to 12k,8k.

03:16 Slewing to WT10_245 for initial collimation pointing/copointing.

03:17 Running dofpia, /first on the field of WT10_245.

03:19 Sky brightness is 17.48 (keeping an eye on it, since we may start LBC observations before 18 deg twilight and I'll note sky brightness when we start).

03:30 added +2000 nm Z8 to DX, -1500nm Z7 to SX. Offset because field of WT10_245 had few pupils.

03:30 dofpia,/first,/x2. sky 19.38. dofpia is adding a lot of spherical, and blue z4 is oscillating. Shift S to stop and go back to WT10_245.

03:39 WT10_245 observation - copoint.

03:45 Slew to field of OB20. Sky background 21.07. Offset by 15 arcmin E to focus/collimate. Seeing is not so great... 1.4" estimated for Red channel and >2" for Blue channel.

03:53 OB20 started. (Still in twilight --- sky brightness 21.60). Guider FWHM values are not as bad as FPIA estimates: 1.1" (B) and 1.2" (R). DIMM says 0.7". Who to believe?!
  • 035421 (b) and 035415 (r) are the first pair of images: ~1" seeing, but backgrounds are a a little high, partly due to Saturn. 12k on blue and 18k on red...
  • blue background going down to 10k, but red is still 18k.
  • 040455 (b) and 040448 (r) 1.1". Red background is still 18k at 040929 (sky background 21.93).

04:17 Still have about 18k counts in center of Red. Going to change exptime to 160 sec, for both blue and red. This means we'll repeat the OB 1.5 times (not 2x but 3x).

04:20 Go to offset field to run dofpia again (indeed it is correcting some coma we saw at the end of the last sequence - hope in the right sense). Sky brightness 21.95 now.
  • blue side taking a while to converge, and est seeing is much greater than for red side (1.38 vs 0.8). Something funny going on with pupils... but converged in 5.

04:28 Started OB20edit (using 160 sec exposure times for Red background <~ 12000 cnts).
  • FWHM on blue (42844) is 5.4 pix(!) but on red (42840) is 3.6 pix (0.8"). Guider reports 1.2, 1.1" FWHM. From 2nd pair, see that blue IQ is not good...

04:46 Focus/collimate on offset field again. Now blue is subtracting some of the z11/z22 added at the start of the night and that red had already started to remove on the previous dofpia.

04:52 Send OB20edit (160s) again (2nd time).
  • 045320 (b) and 045317 (r) about 5 pix (b) and 3.8 pix (r).
  • DIMM seeing 1.1".
  • A top to bottom focus gradient on both blue and red, with best image quality at center and towards top.

05:11 Starting 3rd OB20edit.

05:28 Reconfigure to MODS1.

05:50
  • Collimating on-axis.
  • acqMODS -p PDS456clumps.acq
  • ask to change filter to B_Bessel (no, guide wavelength is not changed in TCS).
    • so, used B_Bessel to guide and acquired at Blue.
  • acqMODS -i PDS46clumps.acq
  • acqMODS -a PDS456clumps.acq (confirmatory thru-slit image is 0038)
  • execMODS PDS456_600sec.obs (will take 2 dithered 600 sec exposures and a 180 sec quasar exposure).
    • guider seeing is 0.9" (B_Bessel).
  • offset back to quasar "offset 2.035 1.42 rel". 180 sec exposure on quasar completed at 06:35. images 41 and 49
  • g_sdss and i_sdss 15-sec x 4 images

06:41
  • acqMODS Q1614+051 PA=35 ... took a long time to collimate - 2 field images taken. thru-slit image is 0057. offset by dx=-0.185 (0058)
  • 06:54 execMODS Q1614+051 will do 3 600 sec exposures.
  • Seeing is 0.9" on guider.

07:31 acqMODS bd33d2642.acq (problem with too many ds9s? modsAlign complained...)
  • 07:39 execMODS bd33d2642.obs
  • and through 1" slit (5" slit and 1" slit positions compare well, see 63 vs 58) --- will run script without checking centering).
  • 07:52 execMODS bd33d2642_1sec.obs (2 x 120s exps)

08:02 acqMODS Q1702+298
  • execMODS Q1702+298 3x600 (this one actually seems fainter than Q1614, i_sdss=19.6 vs 19.22).

08:47 acqMODS V339del at PA=-27 deg (HA=-46 deg now)
  • object saturated in just the 10-sec. Centering was not good. Adjusted by -0.4" and took another thru-slit image (78), this one 7-sec.
  • need yet another 0.4" offset to the left. (79)
  • another (0.2"). Looks good (80).
    • Strange that modsAlign did not work to center it well. The star was saturated in the field image, but the "a" seemed to give a good x,y center.
  • peak counts about 44000 --- looks OK. Interesting that the peak counts seemed to diminish after the first, but maybe just seeing variations. Timescale of object?
  • Unfortunately, we cannot repeat the script. We should switch to LBCs.

09:20 Switching to LBCs.
  • 09:36 Focus/collimate on OB21 field. Blue pupils have a lot of coma, and we have to give -1500nm of Z7 (this is removing what we had added at the start of the night).
  • 09:46 Running OB21
    • FWHM from LBC tech chip guide images is 1.0"
    • 3.7 pix FWHM on blue and 3.2 pix on the red... nice! and blue IQ is better than at start of night.
  • 09:56 Focus/collimate on OB22 field. Converged in two iterations.
  • 10:02 Running OB22
    • from the start, the tech chip images show 1.3" arcsec. But DIMM reads 0.8" still and FPIA estimated seeing values were 0.6 (red) and 0.8 (blue).
    • sky brightness 21.52 at 10:07.
    • first images: 100255 (red) has FWHM ~4 (0.9") and round; 100259 (blue) has FWHM ~5.1 (1.2") and elongated, increasing on 2nd pair. 6.9 (1.55") and 5.8 (1.3").
      • DIMM reads 1.35". Indeed it spiked up about 5 minutes ago.
    • Focus/collimate Sky brightness 21.54. FPIA converged in 1 iteration --- really no need.
    • 10:30 Start OB22 again. Sky brightness 21.54. Blue images still elongated, red appear slightly elongated, but OK. FWHM ~ 6.2 (blue) and ~5.3 (red)
      • 10:45 sky background is 21.47. On 104532 (b), 104514 (r). Finished at 10:54 with sky background 21.23.
    • Sky background on images is not too high, so we'll repeat a 3rd time. Unfortunately, image quality is not good - seeing in this direction and elevation (40 deg) is >1"
      • images 105534 and 105023 have backgrounds ~12000 counts. Sky brightness now 20.67 (11:02 UT). Will stop after next exposure.
    • Stopping at 11:09 sky brightness is 19.89, and last red exposure had 20,000 counts.
  • 11:35 Starting twilight flat fields at g' and r'... PA=0 up to 30,000 in g' and 13,000 counts in r'.
    • PA=180 g' and r'. Took 1-2 before stopping because g' counts were high. r' was up to 30,000.
    • PA=0 S and R flats. Blue rotator did not move and no blue flats were being taken. Stopped & reran the OB... rotator is moving now. But, too late, blue is saturated.

11:51 Closing & going to zenith.

12:06 Starting MODS and LBC calibrations. Telescope @ zenith and lights off in chamber.
  • grpixflats.cal
  • LBC biases
  • grlamps.cal
  • execMODS -b LS60x1.0 slitflats.cal
  • execMODS -b LS60x0.8 slitflats.cal
  • execMODS imflats_edit.cal
    • edited exptimes/lamp levels/lamps and readout region of interest from standard script.
    • z flats have very different levels in each quadrant: Q1, Q2, Q3, Q4 are 14k, 21k, 5k, 11k... something to do with the roi? or z?
  • execMODS bias1K.cal
  • test red channel: i flat looks OK, z flat again has very different values in each quadrant. Increase exptime by x2, but no change. Decrease lamp level and exptime to 1.5/3-sec; they had been saturated.
    • retake the z flats
  • execMODS bias8K.cal

14:10
  • Putting MODS to sleep
  • LBCs are off.
  • LUCI1 MOS remains unresponsive (IT #5223).

-- %USERSIG{OlgaKuhn - 2014-06-23}%

Comments

 
Topic revision: r6 - 24 Jun 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback